# Direct and Converse Theorems in Estimating Signals with Sublinear Sparsity

### v1

@@ Appendix E Proof of Lemma 3
In (158), $\boldsymbol{a}\in\mathbb{R}^{N}$ has independent standard Gaussian elements while the noise $\omega_{N+1}$ is independent of $\{\boldsymbol{x},\boldsymbol{y},\boldsymbol{a}\}$. The coefficient $\alpha>0$ has been introduced to control SNR in the perturbation channel. The limit $\alpha\to 0$ is considered to eliminate the influence of the perturbation channel from the final result. The overall channel is represented with $\tilde{\boldsymbol{y}}=[\boldsymbol{y}^{\mathrm{T}},y_{N+1}]^{\mathrm{T}}$.
We first connect $I(\boldsymbol{x};\tilde{\boldsymbol{y}}|\boldsymbol{a})$ with the square error. Consider the posterior mean estimator $\hat{\boldsymbol{x}}_{\mathrm{opt}}(\boldsymbol{y})=\mathbb{E}[\boldsymbol{x}|\boldsymbol{y}]$ of $\boldsymbol{x}$ given $\boldsymbol{y}$. Since $\{\hat{\boldsymbol{x}}_{\mathrm{opt}}(\boldsymbol{y}),y_{N+1}\}$ is a sufficient statistic in estimating $\boldsymbol{x}$ from $\tilde{\boldsymbol{y}}$ conditioned on $\boldsymbol{a}$, we have $I(\boldsymbol{x},\tilde{\boldsymbol{y}}|\boldsymbol{a})=I(\boldsymbol{x},\hat{\boldsymbol{x}}_{\mathrm{opt}}(\boldsymbol{y}),y_{N+1}|\boldsymbol{a})$. Using the chain rule for mutual information [38], we have
For the first term, we obtain
$I(\boldsymbol{x};\hat{\boldsymbol{x}}_{\mathrm{opt}}(\boldsymbol{y})|\boldsymbol{a})=I(\boldsymbol{x};\hat{\boldsymbol{x}}_{\mathrm{opt}}(\boldsymbol{y}))=I(\boldsymbol{x};\boldsymbol{y})\leq C(p_{\boldsymbol{y}|\boldsymbol{x}}),$ (160)
where $C(p_{\boldsymbol{y}|\boldsymbol{x}})$ denotes the capacity of the channel $p_{\boldsymbol{y}|\boldsymbol{x}}$ under the power constraint $k^{-1}\mathbb{E}[\|\boldsymbol{x}\|_{2}^{2}]\leq|\mathcal{M}|^{-1}\sum_{m\in\mathcal{M}}u_{m}^{2}$.
where the last inequality follows from Jensen’s inequality and $\boldsymbol{a}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I}_{N})$. Using the fact that the posterior mean estimator $\hat{\boldsymbol{x}}_{\mathrm{opt}}(\boldsymbol{y})$ minimizes the mean-square error $\mathbb{E}[\|\Delta\boldsymbol{x}\|_{2}^{2}]$, we arrive at
$-D(p\|q),$ (163)
with $p(\boldsymbol{y})=\mathbb{E}_{\boldsymbol{x}}[p_{\boldsymbol{y}|\boldsymbol{x}}(\boldsymbol{y}|\boldsymbol{x})]$ and $p(\tilde{\boldsymbol{y}}|\boldsymbol{a})=\mathbb{E}_{\boldsymbol{x}}[p(\tilde{\boldsymbol{y}}|\boldsymbol{x},\boldsymbol{a})]$. Combining (159), (162), and (163), we arrive at
with $J(q(\boldsymbol{y}))$ in (28).
We prove the following lower bound on the first term in (164).
Lemma 7
For all $\alpha\in(0,(2u_{\mathrm{max}}^{2})^{-1})$ we have
$+\frac{\alpha}{2k}\mathbb{E}[\|\boldsymbol{x}\|_{2}^{2}]-\log\left(1+f(\alpha)\left\{[1+(N/k-1)^{-1}]^{k}-1\right\}\right),$ (165)
with $f(\alpha)\geq 0$ given in (31).
Proof:
See Appendix H. ∎
We are ready to prove Lemma 3. Applying Lemma 7 to the lower bound (164), we arrive at
$\geq\frac{(1-u_{\mathrm{max}}^{4}\alpha^{2})e^{\frac{\alpha}{k}\mathbb{E}[\|\boldsymbol{x}\|_{2}^{2}]}}{(1+f(\alpha)\{[1+(N/k-1)^{-1}]^{k}-1\})^{2}}e^{-2J(q(\boldsymbol{y}))}-1$ (166)
for all $\alpha\in(0,(2u_{\mathrm{max}}^{2})^{-1})$. Thus, Lemma 3 holds.
Appendix F Proof of Lemma 5
@@
Appendix G Proof of Lemma 8
@@ Appendix H Proof of Lemma 7
$\lambda_{N+1}=1+\frac{\alpha}{k}\boldsymbol{a}^{\mathrm{T}}\mathbb{E}\left[\boldsymbol{x}\boldsymbol{x}^{\mathrm{T}}\right]\boldsymbol{a}.$ (177)
In particular, we use $\boldsymbol{a}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I}_{N})$ to have
@@
Using $p(\tilde{\boldsymbol{y}}|\boldsymbol{a})=\mathbb{E}_{\boldsymbol{x}}[p_{\boldsymbol{y}|\boldsymbol{x}}(\boldsymbol{y}|\boldsymbol{x})p(y_{N+1}|\boldsymbol{x},\boldsymbol{a})]$ and the definition of $y_{N+1}$ in (158) yields
@@
We repeat the derivation of (94): For notational simplicity, without loss of generality, we assume the true support set $\mathcal{S}_{\boldsymbol{x}}=\{1,\ldots,k\}$. Using the representation of $\mathcal{X}_{k}^{N}(\mathcal{U})$ in (13) and decomposing $\boldsymbol{x}^{\prime}\in\mathcal{T}_{w,\boldsymbol{m}}^{N}(\mathcal{S}_{\boldsymbol{x}})$ into $\boldsymbol{x}_{1}^{\prime}\in\mathcal{T}_{k-w,\boldsymbol{m}_{1}}^{k,1}$ and $\boldsymbol{x}_{2}^{\prime}\in\mathcal{T}_{w,\boldsymbol{m}_{2}}^{N-k,2}$ in (48) and (49), we have
@@
$\bar{\chi}_{0,\boldsymbol{m}}=p_{\boldsymbol{y}|\boldsymbol{x}}\left(\boldsymbol{y}\left|\begin{bmatrix}\boldsymbol{u}_{\boldsymbol{m}}\\ \boldsymbol{0}\end{bmatrix}\right.\right)\mathbb{E}\left[\left.e^{(\alpha/k)\boldsymbol{a}_{1}^{\mathrm{T}}\boldsymbol{u}_{\boldsymbol{m}}\boldsymbol{a}_{1}^{\mathrm{T}}\boldsymbol{x}_{1}}\right|\boldsymbol{x}\right],$ (183)
$\bar{\chi}_{k,\boldsymbol{m}}=\sum_{\boldsymbol{x}_{2}^{\prime}\in\mathcal{T}_{k,\boldsymbol{m}}^{N-k,2}}p_{\boldsymbol{y}|\boldsymbol{x}}\left(\boldsymbol{y}\left|\begin{bmatrix}\boldsymbol{0}\\ \boldsymbol{x}_{2}^{\prime}\end{bmatrix}\right.\right)\mathbb{E}\left[\left.e^{(\alpha/k)\boldsymbol{a}_{2}^{\mathrm{T}}\boldsymbol{x}_{2}^{\prime}\boldsymbol{a}_{1}^{\mathrm{T}}\boldsymbol{x}_{1}}\right|\boldsymbol{x}\right],$ (184)
@@
for $w\in\{1,\ldots,k-1\}$, with $\boldsymbol{a}_{1}=\boldsymbol{a}_{\mathcal{S}_{\boldsymbol{x}}}$, $\boldsymbol{a}_{2}=\boldsymbol{a}_{\mathcal{N}\setminus\mathcal{S}_{\boldsymbol{x}}}$, $\boldsymbol{x}_{1}=\boldsymbol{x}_{\mathcal{S}_{\boldsymbol{x}}}$, $\boldsymbol{x}_{1}^{\prime}=\boldsymbol{x}_{\mathcal{S}_{\boldsymbol{x}}}^{\prime}$, $\boldsymbol{x}_{2}^{\prime}=\boldsymbol{x}_{\mathcal{N}\setminus\mathcal{S}_{\boldsymbol{x}}}^{\prime}$, and
@@
For $w=0$ we use Proposition 6 with $\boldsymbol{u}_{1}=\sqrt{\alpha/k}\boldsymbol{u}_{\boldsymbol{m}}$, $\boldsymbol{u}_{2}=\sqrt{\alpha/k}\boldsymbol{x}_{1}$, and $v=0$ to have
$\mathbb{E}\left[\left.e^{(\alpha/k)\boldsymbol{a}_{1}^{\mathrm{T}}\boldsymbol{u}_{\boldsymbol{m}}\boldsymbol{a}_{1}^{\mathrm{T}}\boldsymbol{x}_{1}}\right|\boldsymbol{x}\right]$
$=\left\{[1-(\alpha/k)\boldsymbol{u}_{\boldsymbol{m}}^{\mathrm{T}}\boldsymbol{x}_{1}]^{2}-(\alpha/k)^{2}\|\boldsymbol{u}_{\boldsymbol{m}}\|_{2}^{2}\|\boldsymbol{x}_{1}\|_{2}^{2}\right\}^{-1/2}$
@@
For $w=k$ we evaluate the expectation over $\boldsymbol{a}_{2}$ and subsequently over $\boldsymbol{a}_{1}$ to obtain
$\mathbb{E}\left[\left.e^{(\alpha/k)\boldsymbol{a}_{2}^{\mathrm{T}}\boldsymbol{x}_{2}^{\prime}\boldsymbol{a}_{1}^{\mathrm{T}}\boldsymbol{x}_{1}}\right|\boldsymbol{x}\right]=\mathbb{E}\left[\left.e^{\frac{\alpha^{2}}{2k^{2}}\|\boldsymbol{u}_{\boldsymbol{m}}\|_{2}^{2}(\boldsymbol{a}_{1}^{\mathrm{T}}\boldsymbol{x}_{1})^{2}}\right|\boldsymbol{x}\right]$
@@
$\bar{\chi}(\boldsymbol{y},\boldsymbol{x})$ $\leq(1-2u_{\mathrm{max}}^{2}\alpha)^{-1/2}\Delta p(\boldsymbol{y})+\frac{(1-u_{\mathrm{max}}^{4}\alpha^{2})^{-1/2}}{|\mathcal{X}_{k}^{N}(\mathcal{U})|}$
$\cdot\sum_{\boldsymbol{m}\in\mathcal{M}^{k}}\sum_{\boldsymbol{x}_{1}^{\prime}\in\mathcal{T}_{0,\boldsymbol{m}_{1}}^{k,1}}\sum_{\boldsymbol{x}_{2}^{\prime}\in\mathcal{T}_{k,\boldsymbol{m}}^{N-k,2}}p_{\boldsymbol{y}|\boldsymbol{x}}(\boldsymbol{y}|\boldsymbol{x}^{\prime}),$ (194)
@@
$\Delta p(\boldsymbol{y})=\sum_{w=0}^{k-1}\sum_{\boldsymbol{m}\in\mathcal{M}^{k}}\sum_{\boldsymbol{x}_{1}^{\prime}\in\mathcal{T}_{k-w,\boldsymbol{m}_{1}}^{k,1}}\sum_{\boldsymbol{x}_{2}^{\prime}\in\mathcal{T}_{w,\boldsymbol{m}_{2}}^{N-k,2}}\frac{p_{\boldsymbol{y}|\boldsymbol{x}}(\boldsymbol{y}|\boldsymbol{x}^{\prime})}{|\mathcal{X}_{k}^{N}(\mathcal{U})|}.$ (195)
We are ready to prove Lemma 7. Applying $\mathbb{E}[\log\chi(\tilde{\boldsymbol{y}},\boldsymbol{a})]\leq\mathbb{E}[\log\bar{\chi}(\boldsymbol{y},\boldsymbol{x})$ and (194) to (179), we obtain
$\mathbb{E}\left[\log\frac{p(\boldsymbol{y})p(y_{N+1}|\boldsymbol{x},\boldsymbol{a})}{p(\tilde{\boldsymbol{y}}|\boldsymbol{a})}\right]\geq-\mathbb{E}\left[\log\left\{1+f(\alpha)\frac{\Delta p(\boldsymbol{y})}{p(\boldsymbol{y})}\right\}\right]$
$+\frac{1}{2}\log\left(1-u_{\mathrm{max}}^{4}\alpha^{2}\right)+\frac{\alpha}{2k}\mathbb{E}[\|\boldsymbol{x}\|_{2}^{2}]$ (196)
for all $\alpha\in(0,(2u_{\mathrm{max}}^{2})^{-1})$, with $f(\alpha)\geq 0$ given in (31). Using Jensen’s inequality for the first term in (196) yields
$\mathbb{E}\left[\log\left\{1+f(\alpha)\frac{\Delta p(\boldsymbol{y})}{p(\boldsymbol{y})}\right\}\right]\leq\log\left(1+f(\alpha)\mathbb{E}\left[\frac{\Delta p(\boldsymbol{y})}{p(\boldsymbol{y})}\right]\right)$
$=\log\left(1+f(\alpha)\int_{\mathbb{R}^{N}}\Delta p(\boldsymbol{y})d\boldsymbol{y}\right).$ (197)
From $|\mathcal{X}_{k}^{N}(\mathcal{U})|=|\mathcal{M}|^{k}\binom{N}{k}$ we utilize the definition of $\Delta p(\boldsymbol{y})$ in (195) and the normalization of $p_{\boldsymbol{y}|\boldsymbol{x}}$ over $\boldsymbol{y}$ to have
$\int_{\mathbb{R}^{N}}\Delta p(\boldsymbol{y})d\boldsymbol{y}$ $=\sum_{w=0}^{k-1}\binom{k}{w}\binom{N-k}{w}\binom{N}{k}^{-1}$
$\leq\{1+(N/k-1)^{-1}\}^{k}-1,$ (198)
where the last inequality follows from the upper bound (100) and the binomial theorem. Combining these results, we obtain

### v3

@@ Appendix E Proof of Lemma 5
Appendix F Proof of Lemma 7
@@
Appendix G Proof of Lemma 3
In (173), $\boldsymbol{a}\in\mathbb{R}^{N}$ has independent standard Gaussian elements while the noise $\omega_{N+1}$ is independent of $\{\boldsymbol{x},\boldsymbol{y},\boldsymbol{a}\}$. We recall that $\boldsymbol{x}$ is uniformly distributed on $\mathcal{X}_{k}^{N}(\mathcal{U})$, while $\boldsymbol{y}$ is defined in the AWGN channel (1). The coefficient $\alpha>0$ has been introduced to control SNR in the perturbation channel. The limit $\alpha\to 0$ is considered to eliminate the influence of the perturbation channel from the final result. The overall channel is represented with $\tilde{\boldsymbol{y}}=[\boldsymbol{y}^{\mathrm{T}},y_{N+1}]^{\mathrm{T}}$.
We first connect $I(\boldsymbol{x};\tilde{\boldsymbol{y}}|\boldsymbol{a})$ with the square error. Consider the posterior mean estimator $\hat{\boldsymbol{x}}_{\mathrm{opt}}(\boldsymbol{y})=\mathbb{E}[\boldsymbol{x}|\boldsymbol{y}]$ of $\boldsymbol{x}$ given $\boldsymbol{y}$. Since $\{\hat{\boldsymbol{x}}_{\mathrm{opt}}(\boldsymbol{y}),y_{N+1}\}$ is a sufficient statistic in estimating $\boldsymbol{x}$ from $\tilde{\boldsymbol{y}}$ conditioned on $\boldsymbol{a}$, we have $I(\boldsymbol{x};\tilde{\boldsymbol{y}}|\boldsymbol{a})=I(\boldsymbol{x};\hat{\boldsymbol{x}}_{\mathrm{opt}}(\boldsymbol{y}),y_{N+1}|\boldsymbol{a})$. Using the chain rule for mutual information [39], we have
For the first term, we obtain
$I(\boldsymbol{x};\hat{\boldsymbol{x}}_{\mathrm{opt}}(\boldsymbol{y})|\boldsymbol{a})=I(\boldsymbol{x};\hat{\boldsymbol{x}}_{\mathrm{opt}}(\boldsymbol{y}))=I(\boldsymbol{x};\boldsymbol{y})\leq C_{\mathrm{AWGN}},$ (175)
where $C_{\mathrm{AWGN}}$ denotes the capacity (28) of the AWGN channel (1).
where the last inequality follows from Jensen’s inequality and $\boldsymbol{a}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I}_{N})$. Using the fact that the posterior mean estimator $\hat{\boldsymbol{x}}_{\mathrm{opt}}(\boldsymbol{y})$ minimizes the mean-square error, we arrive at
$-D(p\|q),$ (178)
with $p(\boldsymbol{y})=\mathbb{E}_{\boldsymbol{x}}[p_{\boldsymbol{y}|\boldsymbol{x}}(\boldsymbol{y}|\boldsymbol{x})]$ and $p(\tilde{\boldsymbol{y}}|\boldsymbol{a})=\mathbb{E}_{\boldsymbol{x}}[p(\tilde{\boldsymbol{y}}|\boldsymbol{x},\boldsymbol{a})]$. Combining (174), (175), (177), and (178), we arrive at
with $J(q(\boldsymbol{y}))$ in (III-C).
We prove the following lower bound on the first term in (179).
Lemma 10
Suppose that $p_{\boldsymbol{y}|\boldsymbol{x}}$ is permutation-invariant, i.e. $p_{\boldsymbol{y}|\boldsymbol{x}}(\boldsymbol{P}\boldsymbol{y}|\boldsymbol{P}\boldsymbol{x})=p_{\boldsymbol{y}|\boldsymbol{x}}(\boldsymbol{y}|\boldsymbol{x})$ for any permutation matrix $\boldsymbol{P}$. Then, for all $\alpha\in(0,(2u_{\mathrm{max}}^{2})^{-1})$ we have
$+\frac{\alpha}{2k}\mathbb{E}[\|\boldsymbol{x}\|_{2}^{2}]-\log\left\{1+f(\alpha)\left[(1-k/N)^{-k}-1\right]\right\},$ (180)
with $f(\alpha)\geq 0$ given in (32).
Proof:
See Appendix H. ∎
We are ready to prove Lemma 3. Since the AWGN channel (1) implies the permutation invariance of $p_{\boldsymbol{y}|\boldsymbol{x}}$, we can apply Lemma 10 to the lower bound (179),
$\geq\frac{(1-u_{\mathrm{max}}^{4}\alpha^{2})e^{\frac{\alpha}{k}\mathbb{E}[\|\boldsymbol{x}\|_{2}^{2}]}}{\{1+f(\alpha)[(1-k/N)^{-k}-1]\}^{2}}e^{-2J(q(\boldsymbol{y}))}-1$ (181)
for all $\alpha\in(0,(2u_{\mathrm{max}}^{2})^{-1})$. Using $J(q(\boldsymbol{y}))=D(p(\boldsymbol{y})\|q(\boldsymbol{y}))$ in (III-C) for the Gaussian pdf $q(\boldsymbol{y})$, we arrive at Lemma 3.
@@ Appendix H Proof of Lemma 10
$\lambda_{N+1}=1+\frac{\alpha}{k}\boldsymbol{a}^{\mathrm{T}}\mathbb{E}\left[\boldsymbol{x}\boldsymbol{x}^{\mathrm{T}}\right]\boldsymbol{a},$ (182)
where $\boldsymbol{x}$ is uniformly distributed on $\mathcal{X}_{k}^{N}(\mathcal{U})$. In particular, we use $\boldsymbol{a}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I}_{N})$ to have
@@
We recall that $\boldsymbol{y}$ given $\boldsymbol{x}$ is represented with the conditional pdf $p_{\boldsymbol{y}|\boldsymbol{x}}$. Using $p(\tilde{\boldsymbol{y}}|\boldsymbol{a})=\mathbb{E}_{\boldsymbol{x}}[p_{\boldsymbol{y}|\boldsymbol{x}}(\boldsymbol{y}|\boldsymbol{x})p(y_{N+1}|\boldsymbol{x},\boldsymbol{a})]$ and the definition of $y_{N+1}$ in (173) yields
@@
We repeat the derivation of (99): Using the representation of $\mathcal{X}_{k}^{N}(\mathcal{U})$ in (13) and decomposing $\boldsymbol{x}^{\prime}\in\mathcal{T}_{w,\boldsymbol{m}}^{N}(\mathcal{S}_{\boldsymbol{x}})$ into $\boldsymbol{x}_{1}^{\prime}\in\mathcal{T}_{k-w,\boldsymbol{m}_{1}}^{k,1}$ and $\boldsymbol{x}_{2}^{\prime}\in\mathcal{T}_{w,\boldsymbol{m}_{2}}^{N-k,2}$ in (53) and (54), we have
@@
for $w\in\{0,\ldots,k\}$, with $\boldsymbol{a}_{1}=\boldsymbol{a}_{\mathcal{S}_{\boldsymbol{x}}}$, $\boldsymbol{a}_{2}=\boldsymbol{a}_{\mathcal{N}\setminus\mathcal{S}_{\boldsymbol{x}}}$, $\boldsymbol{x}_{1}=\boldsymbol{x}_{\mathcal{S}_{\boldsymbol{x}}}$, $\boldsymbol{x}_{1}^{\prime}=\boldsymbol{x}_{\mathcal{S}_{\boldsymbol{x}}}^{\prime}$, $\boldsymbol{x}_{2}^{\prime}=\boldsymbol{x}_{\mathcal{N}\setminus\mathcal{S}_{\boldsymbol{x}}}^{\prime}$, and
@@
In particular, for $w=0$ we have $\boldsymbol{x}_{1}^{\prime}=\boldsymbol{u}_{\boldsymbol{m}}$ and $\boldsymbol{x}_{2}^{\prime}=\boldsymbol{0}$ while $\boldsymbol{x}_{1}^{\prime}=\boldsymbol{0}$ holds for $w=k$.
@@
For $w=0$, we have $\boldsymbol{x}_{1}^{\prime}=\boldsymbol{u}_{\boldsymbol{m}}$ and $\boldsymbol{x}_{2}^{\prime}=\boldsymbol{0}$ in $e_{0,\boldsymbol{m}}$ given by (189). we use Proposition 6 with $\boldsymbol{u}_{1}=\sqrt{\alpha/k}\boldsymbol{u}_{\boldsymbol{m}}$, $\boldsymbol{u}_{2}=\sqrt{\alpha/k}\boldsymbol{x}_{1}$, and $v=0$ to have
$e_{0,\boldsymbol{m}}$ $=\left\{[1-(\alpha/k)\boldsymbol{u}_{\boldsymbol{m}}^{\mathrm{T}}\boldsymbol{x}_{1}]^{2}-(\alpha/k)^{2}\|\boldsymbol{u}_{\boldsymbol{m}}\|_{2}^{2}\|\boldsymbol{x}_{1}\|_{2}^{2}\right\}^{-1/2}$
@@
For $w=k$, we have $\boldsymbol{x}_{1}^{\prime}=\boldsymbol{0}$ in $e_{k,\boldsymbol{m}}$. Evaluating the expectation over $\boldsymbol{a}_{2}$ and subsequently over $\boldsymbol{a}_{1}$ yields
$e_{k,\boldsymbol{m}}=\mathbb{E}\left[\left.e^{\frac{\alpha^{2}}{2k^{2}}\|\boldsymbol{u}_{\boldsymbol{m}}\|_{2}^{2}(\boldsymbol{a}_{1}^{\mathrm{T}}\boldsymbol{x}_{1})^{2}}\right|\boldsymbol{x}\right]$
@@
$\bar{\chi}(\boldsymbol{y},\boldsymbol{x})\leq$ $(1-2u_{\mathrm{max}}^{2}\alpha)^{-1/2}\left\{p(\boldsymbol{y})-\Delta_{\mathcal{S}_{\boldsymbol{x}}}p(\boldsymbol{y})\right\}$
$+(1-u_{\mathrm{max}}^{4}\alpha^{2})^{-1/2}\Delta_{\mathcal{S}_{\boldsymbol{x}}}p(\boldsymbol{y}),$ (197)
@@
$\Delta_{\mathcal{S}_{\boldsymbol{x}}}p(\boldsymbol{y})=\frac{1}{|\mathcal{X}_{k}^{N}(\mathcal{U})|}\sum_{\boldsymbol{m}\in\mathcal{M}^{k}}\sum_{\boldsymbol{x}_{2}^{\prime}\in\mathcal{T}_{k,\boldsymbol{m}}^{N-k,2}}p_{\boldsymbol{y}|\boldsymbol{x}}(\boldsymbol{y}|\boldsymbol{x}^{\prime}),$ (198)
with $\boldsymbol{x}_{\mathcal{S}_{\boldsymbol{x}}}^{\prime}=\boldsymbol{0}$ and $\boldsymbol{x}_{2}^{\prime}=\boldsymbol{x}_{\mathcal{N}\setminus\mathcal{S}_{\boldsymbol{x}}}^{\prime}$. The quantity $\Delta_{\mathcal{S}_{\boldsymbol{x}}}p(\boldsymbol{y})$ depends on $\mathcal{S}_{\boldsymbol{x}}$ only through the partition of $\boldsymbol{x}^{\prime}$.
We evaluate the quantity (184). Applying (197) and the inequality $\mathbb{E}[\log\chi(\tilde{\boldsymbol{y}},\boldsymbol{a})]\leq\mathbb{E}[\log\bar{\chi}(\boldsymbol{y},\boldsymbol{x})]$ to (184), we obtain